OpenStudy (anonymous):

A bus accelerates at -0.5 m/s2 for 12.5 seconds. If the bus had an initial speed of 31 m/s, what is its new speed? A. 24.75 m/s B. 12.25 m/s C. 6.25 m/s D. 37.25 m/s

OpenStudy (anonymous):

@nermaljean99

OpenStudy (anonymous):

final velocity=initial velocity+(acceleration x time) insert values and thats the answer

OpenStudy (anonymous):

v=31+(0.5x12.5)=37.5?

OpenStudy (anonymous):

37.25 I mean?

OpenStudy (anonymous):

no...close though You did the right substitution but your math is wrong. (-.5m/s^2 times 12.5)=-6.25 then add 31 and the answer is 24.75

OpenStudy (anonymous):

Oh okay! I see now, thank you! :)